Determine the numerical values of γ
γ
for an object moving at speed v=0.01c
v
=
0.01
c
, 0.05c
0.05
c
, 0.10c
0.10
c
, 0.20c
0.20
c
, 0.30c
0.30
c
, 0.40c
0.40
c
, 0.50c
0.50
c
, 0.60c
0.60
c
, 0.70c
0.70
c
, 0.80c
0.80
c
, 0.90c
0.90
c
, and 0.99c
0.99
c
. Make a graph of γ
γ
versus v
v
.
Answer #1

| 7 (1-19 21² = . . ا . هه 2 .ر2- (// = لا 0.01 دوا (1) (1-e-02 ))) (-0.02) = ۱۰ 005 دلار یان.0 = و (1) .27ري. .-1 ) - 1000vi) v = 0.UC, ya bru (1-0.437²2 Y = 1.091 (n) & 0.50, y = (1-6.5 272 8 = 1.154 (VIT) = 0.60, y = (170.678] = 1.25 7 2 (ix) =y VSV 8 7 6 5 >4 3 2 1 0 0 0.2 0.4 0.6 0.8 1 1.2 v(in unit of c)